A project yields an average cash-flow of Rs. 400 lakh with a standard deviation of Rs. 60 lakh. Calculate the following probabilities: (i) Cash flow will be more than Rs. 550 lakh (ii) Cash flow will be between Rs. 450 lakh and Rs. 520 lakh.


1
Expert's answer
2021-07-12T16:47:39-0400

Let "X=" cash flow: "X\sim N(\mu, \sigma^2)."

Given "\mu=400, \sigma=60."

(i)


"P(X>550)=1-P(X\leq 550)"

"=1-P(Z\leq \dfrac{550-400}{60})"

"=1-P(Z\leq 2.5)"

"\approx0.0062"

(ii)


"P(450<X<520)=P(X<520)-P(X\leq 450)"

"=P(Z< \dfrac{520-400}{60})-P(Z\leq \dfrac{450-400}{60})"

"=P(Z<2)-P(Z\leq 0.8333)"

"\approx0.97725-0.79767\approx0.1796"